### v3

# Stochastic Vector Approximate Message Passing with applications to phase retrieval

Abstract
Phase retrieval refers to the problem of recovering a high-dimensional vector $\boldsymbol{x}\in\mathbb{C}^{N}$ from the magnitude of its linear transform $\boldsymbol{z}=A\boldsymbol{x}$, observed through a noisy channel. To improve the ill-posed nature of the inverse problem, it is a common practice to observe the magnitude of linear measurements $\boldsymbol{z}^{(1)}=A^{(1)}\boldsymbol{x},...,\boldsymbol{z}^{(L)}=A^{(L)}\boldsymbol{x}$ using multiple sensing matrices $A^{(1)},...,A^{(L)}$, with ptychographic imaging being a remarkable example of such strategies. Inspired by existing algorithms for ptychographic reconstruction, we introduce stochasticity to Vector Approximate Message Passing (VAMP), a computationally efficient algorithm applicable to a wide range of Bayesian inverse problems. By testing our approach in the setup of phase retrieval, we show the superior convergence speed of the proposed algorithm.
Index Terms—  Phase Retrieval, Bayesian inverse problems, Belief Propagation
1 Introduction
1.1 Background on Phase Retrieval
Phase retrieval is the problem of recovering a high-dimensional vector $\boldsymbol{x}\in\mathbb{C}^{N}$ from the observation $\boldsymbol{y}\in\mathbb{R}^{M}$ modeled by
$\ \boldsymbol{z}=A\boldsymbol{x}\ ,\ y_{\mu}\sim p_{\text{out}}(\ \cdot\ |\ |z_{\mu}|\ )\ \ (\mu=1,...,M)$ (1)
where $A\in\mathbb{C}^{M\times N}$ is the sensing matrix and $p_{\text{out}}(\ \cdot\ |\ |z_{\mu}|\ )$ is a probability measure given the magnitude of the $\mu$-th entry of $\boldsymbol{z}\in\mathbb{C}^{M}$. A typical application that fits this framework is Coherent Diffraction Imaging (CDI) [1], where an object image $\boldsymbol{x}$ is illuminated by coherent light, and the magnitude of diffracted light is observed in the far field.
It is of practical interest how many observations are needed to recover the signal $\boldsymbol{x}$. In recent years, the information-theoretic threshold of sampling ratio has been intensively studied for phase retrieval with large random sensing matrices [2]. From an algorithmic viewpoint, spectral method  [3, 4] is known to be an efficient algorithm for obtaining an estimate $\hat{\boldsymbol{x}}$ correlated with $\boldsymbol{x}$, which is often employed to initialize iterative solvers, including the Wirtinger flow [5], and message passing algorithms [6].
In computational imaging, there are mainly two strategies to improve the ill-posed nature of phase reconstruction – multiple sensing matrices and exploitation of prior knowledge. A remarkable example of the former is ptychography in CDI, where the magnitudes of several linear measurements $\boldsymbol{z}^{(1)}=A^{(1)}\boldsymbol{x},...,\boldsymbol{z}^{(L)}=A^{(L)}\boldsymbol{x}$ are observed: In a typical setting of ptychography, each sensing matrix $A^{(l)}\in\mathbb{C}^{\bar{M}\times N}\ (l=1,...,L)$ is given as $A^{(l)}=FPS^{(l)}$, where $F\in\mathbb{C}^{\bar{M}\times\bar{M}}$ is the 2D DFT, $P\in\mathbb{C}^{\bar{M}\times\bar{M}}$ is a diagonal matrix, and $S^{(l)}\in\{0,1\}^{\bar{M}\times N}$ is a binary matrix consisting of $\bar{M}$ row vectors of the identity matrix $I_{N}$. On the forward model of ptychography, see, for example, [4, 7]. Another example is coded diffraction pattern [8], whose sensing matrices are expressed as $A^{(l)}=FP^{(l)}$, where $F\in\mathbb{C}^{N\times N}$ is the 2D DFT and $P^{(l)}=\text{Diag}(e^{j\theta^{(l)}_{1}},\ldots,e^{j\theta^{(l)}_{N}})$ denotes a random phase modulation. This model corresponds to CDI with ”random phase masks” placed after the object $\boldsymbol{x}$, as illustrated in Fig. 1. Random masks are actively utilized in CDI experiments [9].
Fig. 1: Setup of coded diffraction pattern
Regarding the latter approach, Bayesian inference provides a framework for incorporating prior knowledge, such as sparsity and support constraint, into a statistical model. The Bayes’ theorem for the model (1) is written as
$p(\boldsymbol{x},\boldsymbol{z}|\boldsymbol{y})\propto p_{\text{in}}(\boldsymbol{x})\delta(\boldsymbol{z}-A\boldsymbol{x})\prod_{\mu=1}^{M}p_{\text{out}}(y_{\mu}||z_{\mu}|)$ (2)
wherein the prior $p_{\text{in}}(\boldsymbol{x})$ expresses the prior knowledge on $\boldsymbol{x}$, and $\delta(\cdot)$ is the Dirac delta function. To compute the posterior density $p(\boldsymbol{x},\boldsymbol{z}|\boldsymbol{y})$ in a reasonable computational time, approximations such as the mean-field approach and approximate message passing have been applied to phase retrieval [2, 10, 11]. These techniques can also be combined with deep neural network trained with natural images [12].
1.2 Illustration of the key idea
Our approach is to blend the two techniques stated above: multiple measurements and Bayesian inference. We aim to reconstruct a complex image $\boldsymbol{x}\in\mathbb{C}^{N}$ from $L$ observations $\boldsymbol{y}^{(l)}\in\mathbb{R}^{\bar{M}}$ $(l=1,\ldots,L)$. Using sensing matrices $A^{(1)},\ldots,A^{(L)}\in\mathbb{C}^{\bar{M}\times N}$, the observations are modeled by
$\quad\boldsymbol{z}^{(l)}=A^{(l)}\boldsymbol{x}\ ,\ \boldsymbol{y}^{(l)}\sim p_{\text{out}}(\cdot\mid|\boldsymbol{z}^{(l)}|)\quad(l=1,...,L)$ (3)
with $p_{\text{out}}(\boldsymbol{y}\mid|\boldsymbol{z}|)\triangleq\prod_{\mu}p_{\text{out}}(y_{\mu}\mid|z_{\mu}|)$. By defining the concatenated sensing matrix and observation as
$A=\begin{pmatrix}A^{(1)}\\ \vdots\\ A^{(L)}\end{pmatrix},\ \boldsymbol{z}=\begin{pmatrix}\boldsymbol{z}^{(1)}\\ \vdots\\ \boldsymbol{z}^{(L)}\end{pmatrix},\ \boldsymbol{y}=\begin{pmatrix}\boldsymbol{y}^{(1)}\\ \vdots\\ \boldsymbol{y}^{(L)}\end{pmatrix}$ (4)
the model (3) is reduced to model (1). However, crucial to our algorithm is sequential access to the observations $\boldsymbol{y}^{(1)},\ldots,\boldsymbol{y}^{(L)}$ which leads to faster convergence, compared to accessing all of the observations at once. Indeed, sequential ptychographic solvers such as ePIE [13] is known to effectively avoid bad local minima, as detailed in  [14].
To incorporate multiple measurements into a Bayesian framework, we extend Vector Approximate Message Passing (VAMP)  [15, 16], an efficient inference algorithm for generalized linear models. Our extension, coined Stochastic VAMP, allows ”mini-batch learning” commonly employed in ptychographic reconstruction[14] and speeds up convergence. This method is analogous to Stochastic Gradient Descent in non-convex optimization, where gradients are computed from data subsets, hence the name Stochastic VAMP.
2 Stochastic VAMP
2.1 Derivation of VAMP
We review the derivation of VAMP based on expectation propagation (EP)  [15, 16]. For the Bayesian problem in Eq. (1), we start with an equivalent model:
$p(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\boldsymbol{z}_{1},\boldsymbol{z}_{2}|\boldsymbol{y})\propto$ $p_{\text{in}}(\boldsymbol{x}_{1})\delta(\boldsymbol{x}_{1}-\boldsymbol{x}_{2})\times$ (5)
$\delta(A\boldsymbol{x}_{2}-\boldsymbol{z}_{2})\delta(\boldsymbol{z}_{1}-\boldsymbol{z}_{2})p_{\text{out}}(\boldsymbol{y}||\boldsymbol{z}_{1}|)$
The corresponding factor graph is shown in Fig.2 (a). In belief propagation (BP) [17], probability distributions called messages are passed through the graph to form beliefs, which are the estimates of marginal distributions of each variable. EP approximates messages by using Gaussian. Applying EP to the graph in Fig.2 (a) yields the VAMP algorithm.
Fig. 2: (a) Factor graph (b) Denoising (c) LMMSE estimation
Here, we outline the update rules of VAMP. Each iteration includes two steps: denoising and LMMSE estimation. In the denoising part, $\boldsymbol{x}_{1}$ node inputs the message $\mathcal{N}(\boldsymbol{x}_{1}|\boldsymbol{r}_{1},\gamma_{1}^{-1}I_{N})$ from $\delta(\boldsymbol{x}_{1}-\boldsymbol{x}_{2})$ node. [^1] The density of the complex circular Gaussian distribution is given as $\mathcal{N}(\boldsymbol{x}|\boldsymbol{r},\gamma^{-1}I_{N})\propto\exp(-\gamma||\boldsymbol{x}-\boldsymbol{r}||_{2}^{2})$, wherein $\boldsymbol{x},\boldsymbol{r}\in\mathbb{C}^{N}$ and $\gamma>0$. This creates the belief $b(\boldsymbol{x}_{1})\propto p_{\text{in}}(\boldsymbol{x}_{1})\mathcal{N}(\boldsymbol{x}_{1}|\boldsymbol{r}_{1},\gamma_{1}^{-1}I_{N})$. Denoting expectation w.r.t. measure $b(\boldsymbol{x}_{1})$ by $\mathbb{E}[\ \cdot\ |b(\boldsymbol{x}_{1})]$, the denoiser functions are defined as
$\boldsymbol{g}_{x}(\boldsymbol{r}_{1},\gamma_{1})$ $=\mathbb{E}[\boldsymbol{x}_{1}|b(\boldsymbol{x}_{1})]$ (6)
$\boldsymbol{g}^{\prime}_{x}(\boldsymbol{r}_{1},\gamma_{1})$ $=\frac{1}{N}\mathbb{E}\left[||\boldsymbol{x}_{1}-\boldsymbol{g}_{x}(\boldsymbol{r}_{1},\gamma_{1})||_{2}^{2}|b(\boldsymbol{x}_{1})\right]$ (7)
$b(\boldsymbol{x}_{1})$ is approximated by $\mathcal{N}(\boldsymbol{x}_{1}|\hat{\boldsymbol{x}}_{1},\eta_{1}^{-1}I_{N})$ with
$\hat{\boldsymbol{x}}_{1}=\boldsymbol{g}_{x}(\boldsymbol{r}_{1},\gamma_{1})\ ,\ \eta_{1}^{-1}=\boldsymbol{g}^{\prime}_{x}(\boldsymbol{r}_{1},\gamma_{1})$ (8)
Then, the message $\mathcal{N}(\boldsymbol{x}_{2}|\boldsymbol{r}_{2},\gamma_{2}^{-1}I_{N})$ is sent to $\boldsymbol{x}_{2}$ node from $\delta(\boldsymbol{x}_{1}-\boldsymbol{x}_{2})$ node, which, by the rules of BP, is given by
$\gamma_{2}=\eta_{1}-\gamma_{1}\ ,\ \boldsymbol{r}_{2}=(\eta_{1}\hat{\boldsymbol{x}}_{1}-\gamma_{1}\boldsymbol{r}_{1})/\gamma_{2}$ (9)
Likewise, the message $\mathcal{N}(\boldsymbol{z}_{1}|\boldsymbol{p}_{1},\tau_{1}^{-1}I_{M})$ is input to $\boldsymbol{z}_{1}$ node. $\hat{\boldsymbol{z}}_{1},\lambda_{1},\boldsymbol{p}_{2}$, and $\tau_{2}$ are give by
$\boldsymbol{g}_{z}(\boldsymbol{p}_{1},\tau_{1})=\mathbb{E}[\boldsymbol{z}_{1}|b(\boldsymbol{z}_{1})]$ (10)
$\boldsymbol{g}^{\prime}_{z}(\boldsymbol{p}_{1},\tau_{1})=\frac{1}{M}\mathbb{E}\left[||\boldsymbol{z}_{1}-\boldsymbol{g}_{z}(\boldsymbol{p}_{1},\tau_{1})||_{2}^{2}|b(\boldsymbol{z}_{1})\right]$ (11)
$\hat{\boldsymbol{z}}_{1}=\boldsymbol{g}_{z}(\boldsymbol{p}_{1},\tau_{1})\ ,\ \lambda_{1}^{-1}=\boldsymbol{g}^{\prime}_{z}(\boldsymbol{p}_{1},\tau_{1})$ (12)
$\tau_{2}=\lambda_{1}-\tau_{1}\ ,\ \boldsymbol{p}_{2}=(\lambda_{1}\hat{\boldsymbol{z}}_{1}-\tau_{1}\boldsymbol{p}_{1})/\tau_{2}$ (13)
with $b(\boldsymbol{z}_{1})\propto\mathcal{N}(\boldsymbol{z}_{1}|\boldsymbol{p}_{1},\tau_{1}^{-1}I_{M})p_{\text{out}}(\boldsymbol{y}||\boldsymbol{z}_{1}|)$.
In the LMMSE estimation, joint belief on $\boldsymbol{x}_{2}$ and $\boldsymbol{z}_{2}$ is given as $b(\boldsymbol{x}_{2},\boldsymbol{z}_{2})\propto\mathcal{N}(\boldsymbol{x}_{2}|\boldsymbol{r}_{2},\gamma_{2}^{-1}I_{N})\times\delta(A\boldsymbol{x}_{2}-\boldsymbol{z}_{2})\mathcal{N}(\boldsymbol{z}_{2}|\boldsymbol{p}_{2},\tau_{2}^{-1}I_{M})$. Marginalizing $b(\boldsymbol{x}_{2},\boldsymbol{z}_{2})$ w.r.t. $\boldsymbol{z}_{2}$, we have $b(\boldsymbol{x}_{2})\propto\mathcal{N}(\boldsymbol{x}_{2}|\boldsymbol{r}_{2},\gamma_{2}^{-1}I_{N})\times\mathcal{N}(A\boldsymbol{x}_{2}|\boldsymbol{p}_{2},\tau_{2}^{-1}I_{M})$. This is then approximated by $\mathcal{N}(\boldsymbol{x}_{2}|\hat{\boldsymbol{x}}_{2},\eta_{2}^{-1}I_{N})$ with
$\hat{\boldsymbol{x}}_{2}=Q(\gamma_{2}\boldsymbol{r}_{2}+\tau_{2}A^{H}\boldsymbol{p}_{2})\ ,\ \eta_{2}^{-1}=\frac{1}{N}\text{Tr}(Q)$ (14)
$Q=(\gamma_{2}I_{N}+\tau_{2}A^{H}A)^{-1}$ (15)
In the rest of this paper, we focus on the case where $A^{H}A$ is diagonal, thus the matrix inversion in Eq. (15) is not the dominant computational cost. For general sensing matrices, per-iteration matrix inversion can be avoided by pre-computing the singular value decomposition of $A$. Similarly, $b(\boldsymbol{z}_{2})$ is approximated by $\mathcal{N}(\boldsymbol{z}_{2}|\hat{\boldsymbol{z}_{2}},\lambda_{2}^{-1}I_{M})$ with
$\hat{\boldsymbol{z}}_{2}=A\hat{\boldsymbol{x}}_{2}\ ,\lambda_{2}^{-1}=\frac{1}{M}\text{Tr}(AQA^{H})$ (16)
The messages $\mathcal{N}(\boldsymbol{x}_{1}|\boldsymbol{r}_{1},\gamma_{1}^{-1}I_{N})$ and $\mathcal{N}(\boldsymbol{z}_{1}|\boldsymbol{p},\tau_{1}^{-1}I_{M})$ are updated according to
$\gamma_{1}=\eta_{2}-\gamma_{2}\ ,\boldsymbol{r}_{1}=(\eta_{2}\hat{\boldsymbol{x}}_{2}-\gamma_{2}\boldsymbol{r}_{2})/\gamma_{1}$ (17)
$\tau_{1}=\lambda_{2}-\tau_{2}\ ,\boldsymbol{p}_{1}=(\lambda_{2}\hat{\boldsymbol{z}}_{2}-\tau_{2}\boldsymbol{p}_{2})/\tau_{1}$ (18)
which closes the VAMP iteration.
2.2 Stochastic VAMP
A natural extension of VAMP to multiple measurements is derived by EP on the factor graph in Fig.3. Denote the message sent from $\delta(\boldsymbol{z}^{(l)}_{1}-\boldsymbol{z}^{(l)}_{2})$ to $\boldsymbol{z}^{(l)}_{1}$ and the message sent from $\delta(\boldsymbol{z}^{(l)}_{1}-\boldsymbol{z}^{(l)}_{2})$ to $\boldsymbol{z}^{(l)}_{2}$ by $\mathcal{N}\left(\boldsymbol{z}^{(l)}_{1}|\boldsymbol{p}^{(l)}_{1},(\tau^{(l)}_{1})^{-1}I_{M}\right)$ and $\mathcal{N}\left(\boldsymbol{z}^{(l)}_{2}|\boldsymbol{p}^{(l)}_{2},(\tau^{(l)}_{2})^{-1}I_{M}\right)$, respectively. It is straight forward to derive the LMMSE estimator to update $(\boldsymbol{p}^{(l)}_{1},\tau^{(l)}_{1})$ and $(\boldsymbol{p}^{(l)}_{2},\tau^{(l)}_{2})$ for $l=1,...,L$. Note that if we update these $L$ messages in parallel, the resulting algorithm is essentially the same as VAMP with the sensing matrix $A$ and observation $\boldsymbol{y}$ defined in Eq. (4). The only difference from VAMP is that we allow $\{\tau^{(l)}_{1},\tau^{(l)}_{2}\}$ $(l=1,...,L)$ to take different values. However, we can also update $L$ messages sequentially, yelding Stochastic VAMP specified in Algorithm 1.
Fig. 3: Factor graph of stochastic VAMP
Algorithm 1 Stochastic VAMP
1:Sensing matrices $A^{(1)},...,A^{(L)}$, observations $\boldsymbol{y}^{(1)},...,\boldsymbol{y}^{(L)}$, number of iterations $K_{\text{it}}$
2:Initialize $\boldsymbol{r}_{2},\gamma_{2}>0$ and $\boldsymbol{p}_{2},\tau_{2}>0\ \ (l=1,...,L)$
3:for $k=0,1,...,K_{\text{it}}$ do
4:     for $l=1,...,L$ do
5:         //LMMSE estimation
6:         $Q=\left(\gamma_{2}I_{N}+\sum_{l=1}^{L}\tau^{(l)}_{2}(A^{(l)})^{H}A^{(l)}\right)^{-1}$
7:         $\hat{\boldsymbol{x}}_{2}=Q\left(\gamma_{2}\boldsymbol{r}_{2}+\sum_{l=1}^{L}\tau^{(l)}_{2}(A^{(l)})^{H}\boldsymbol{p}^{(l)}_{2}\right)$
8:         $\eta_{2}^{-1}=\frac{1}{N}\text{Tr}(Q)$
9:         $\hat{\boldsymbol{z}}_{2}=A^{(l)}\hat{\boldsymbol{x}}_{2}$
10:         $(\lambda^{(l)}_{2})^{-1}=\frac{1}{M}\text{Tr}\left(A^{(l)}Q(A^{(l)})^{H}\right)$
11:         // Message passing
12:         $\gamma_{1}=\eta_{2}-\gamma_{2}$
13:         $\boldsymbol{r}_{1}=(\eta_{2}\hat{\boldsymbol{x}}_{2}-\gamma_{2}\boldsymbol{r}_{2})/\gamma_{1}$
14:         $\tau_{1}^{(l)}=\lambda_{2}^{(l)}-\tau_{2}^{(l)}$
15:         $\boldsymbol{p}^{(l)}_{1}=(\lambda_{2}^{(l)}\hat{\boldsymbol{z}}_{2}-\tau_{2}^{(l)}\boldsymbol{p}^{(l)}_{2})/\tau_{1}^{(l)}$
16:         //Denoising
17:         $\hat{\boldsymbol{x}}_{1}=\boldsymbol{g}_{x}(\boldsymbol{r}_{1},\gamma_{1})$
18:         $\eta_{1}^{-1}=\boldsymbol{g}^{\prime}_{x}(\boldsymbol{r}_{1},\gamma_{1})$
19:         $\hat{\boldsymbol{z}}^{(l)}_{1}=\boldsymbol{g}_{z}(\boldsymbol{p}^{(l)}_{1},\tau_{1})$
20:         $\lambda_{1}^{-1}=\boldsymbol{g}^{\prime}_{z}(\boldsymbol{p}^{(l)}_{1},\tau_{1})$
21:         // Message passing
22:         $\gamma_{2}=\eta_{1}-\gamma_{1}$
23:         $\boldsymbol{r}_{2}=(\eta_{1}\hat{\boldsymbol{x}}_{1}-\gamma_{1}\boldsymbol{r}_{1})/\gamma_{2}$
24:         $\tau_{2}^{(l)}=\lambda_{1}^{(l)}-\tau_{1}^{(l)}$
25:         $\boldsymbol{p}^{(l)}_{2}=(\lambda_{1}^{(l)}\hat{\boldsymbol{z}}_{1}-\tau_{1}^{(l)}\boldsymbol{p}^{(l)}_{1})/\tau_{2}^{(l)}$
26:     end for
27:end for
28:Return $\hat{\boldsymbol{x}}_{1}$
2.3 Algorithmic detail
It is standard to stabilize the convergence of VAMP by employing the ”damping” method, where $\boldsymbol{r}_{1}$ in Eq. (17) is treated as ”raw” value $\boldsymbol{r}_{\text{raw}}$, and $\boldsymbol{r}_{1}$ is updated as $\boldsymbol{r}_{1}=\rho\boldsymbol{r}_{\text{raw}}+(1-\rho)\boldsymbol{r}_{\text{old}}$, in which $10<\rho<$ 1 and $\boldsymbol{r}_{\text{old}}$ is the previous value of $\boldsymbol{r}_{1}$. Similar technique is applied to the update of $\gamma,\boldsymbol{p},$ and $\tau$. Notably, Stochastic VAMP requires less damping than standard VAMP, contributing to a faster convergence.
When each matrix $A^{(l)}$ have $(A^{(l)})^{H}A^{(l)}$ diagonal, the matrix inversion in line 5 adds only minor complexity to the algorithm. The dominant computational cost comes from the matrix-vector multiplications in lines 6 and 8. In the case of coded diffraction patterns, the cost is $2L$ FFTs per iteration. For general sensing matrices, costly matrix-inversion can still be avoided by using the graphical model introduced for D-VAMP [18] and pre-computing the singular value decompositions of $A^{(1)},...,A^{(L)}$.
In our experiments, we assumed that the SNR and prior are known, but they can be learned from the dataset using methods from  [19]. Although noise learning and multiple restarts can help phase retrieval algorithms avoid bad local minima [10, 11], numerical results indicate that Stochastic VAMP remains stable even without these enhancements. Python codes for the numerical experiments in this paper is available at https://github.com/sacbow/StochasticVAMP.
3 Computational results
3.1 Common setup
In this section, we consider an observation model defined as:
$\boldsymbol{y}=|\boldsymbol{z}+\boldsymbol{w}|$ (19)
where $\boldsymbol{w}$ is the Gaussian noise given as $\boldsymbol{w}\sim\mathcal{N}(\cdot|\boldsymbol{0},\gamma_{w}^{-1}I_{M})$, and ||$\cdot$|| ⋅ | is the element-wise absolute value. The corresponding likelihood function $p_{\text{out}}(\boldsymbol{y}||\boldsymbol{z}|)$ is refered to as the Rician distribution, whose denoising functions are derived in  [11]. We assume a Gaussian prior $p_{\text{in}}(\boldsymbol{x})=\mathcal{N}(\boldsymbol{x}|\boldsymbol{0},I_{N})$, but structured priors can also be used. The total number of observed values is $M=L\bar{M}$, where $L$ is the number of sensing matrices $A^{(1)},...,A^{(L)}$ and $A^{(l)}\in\mathbb{C}^{\bar{M}\times N}$. The sampling ratio is $\alpha\triangleq\frac{M}{N}$. Since the posterior density is invariant w.r.t. global phase shift, the deviation between the true signal $\boldsymbol{x}$ and VAMP estimate $\hat{\boldsymbol{x}}$ is measured by
$\text{NMSE}\triangleq\underset{\theta\in[0,2\pi)}{\operatorname{min}}\frac{||\boldsymbol{x}-e^{j\theta}\hat{\boldsymbol{x}}||^{2}}{||\boldsymbol{x}||^{2}}$ (20)
3.2 Haar distributed sensing matrix
We consider sensing matrices $A^{(l)}$ with $(A^{(l)})^{H}A^{(l)}=I_{N}$ drawn from Haar distribution. In the case of noiseless phase retrieval, the threshold value of $\alpha$ above which VAMP recovers the true signal $\boldsymbol{x}$ is investigated via asymptotic analysis [2], in which the threshold value for Haar-distributed $A$ is estimated as $\alpha_{\text{FR, Alg}}\simeq 2.265$. In our experiment, we used $N=512$, $\alpha=2.4$, and $L=2$, with SNR of 30 dB. We compared Stochastic VAMP with vanilla VAMP, which uses concatenated sensing matrices and observations (Eq.(4)). While VAMP required damping of $\rho=0.9$ for convergence, Stochastic VAMP was stable with less amount of damping ($\rho=0.97$). The true signal $\boldsymbol{x}$ was drawn from the prior, and both algorithms were randomly initialized. The complexity of both algorithms is similar, with $O((L-1)N)$ difference per-iteration. The NMSE is shown against iteration number $k$ in Fig. 4. We observed that the stochasticity significantly accelerates the convergence of the phase retrieval algorithm.
Fig. 4: NMSE versus iteration number for VAMP and Stochastic VAMP.
3.3 Coded diffraction pattern from real world image
In this section, we evaluate the performance of Stochastic VAMP in a more realistic setting. Using the sensing matrix for coded diffraction patterns with $L=3$ random masks, we conducted numerical experiments on a $256\times 256$-pixel real-valued image (Fig.5(a)). The SNR is set to 30 dB. A typical reconstruction is shown in Fig.5(b), which was obtained within 20 iterations with a computational time of only 3.9 seconds. Notably, this rapid convergence is impressive, given that conventional phase retrieval algorithms typically require $10^{2}$ to $10^{3}$ iterations for reliable reconstruction. Although we assumed complex Gaussian prior, the reconstruction can be further improved by using non-negative real prior.
Fig. 5: (a) test image (b) reconstructed image (c) NMSE versus iteration number for Stochastic VAMP.
4 Conclusions
We enhanced VAMP by introducing stochasticity through non-parallel message updates in belief propagation. In phase retrieval with multiple measurements, our approach significantly accelerates convergence while maintaining the per-iteration complexity of conventional algorithms.
Being a Bayesian approach, our method allows for incorporating prior knowledge, such as image sparsity, similar to  [11]. Future work includes applying this approach to ptychographic imaging. Practical ptychographic phase retriecal algorithms such as ePIE perform self-calibration by which both the probe and the object image are reconstructed. Probe reconstruction could be achieved by integrating Stochastic VAMP with techniques from  [20, 21].